Home‎ > ‎Crew 133‎ > ‎

110-ScienceReport

Science Report
01/10/2014
Paula Crock
 
Today was the first day of data collection for the EVA helmet temperature study and the radio 'dead'
zone mapping survey.
 
EVA helmet temps:  Data were collected on two subjects during today's EVA.  The devices seemed to
work properly although one of the internal helmet temperature sensors came out of the helmet during
the EVA so a better placement of the device needs to be found for the old style EVA suits.  I expect
to collect data on two or three crew members per EVA through the rest of the rotation.
 
Radio Dead Zone map:  For the EVA today we headed due east and did regular radio check-ins with the
Hab to see if we could find the line-of-sight distance of the radios (range) and then headed down
into a canyon to check elevation issues with radio communications.   One the way back from the
canyon we collected data on when the isolated hills block radio signal from an otherwise
line-of-sight signal to the Hab.  Coordinates were collected at each call attempt and will be
incorporated with similar tests in future EVAs.  These data will be used to validate the map
developed with the DEM terrain model which is still in development.  The initial test went well and
I plan to use future EVAs to test other terrain features within the 2km walking zone of the Hab.